OpenStudy (anonymous):

Which of the following is the slope between the two points (-1, -1) and (8, 0)? Choose one answer. a. 9 b. -9 c. one ninth d. -one ninth

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

No d

OpenStudy (welshfella):

slope = rise / run = 0-(-1) / 8 - (-1)

OpenStudy (anonymous):

so it is d?

OpenStudy (welshfella):

no

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-1/9

OpenStudy (anonymous):

oh 1/9 i see it

OpenStudy (welshfella):

- double negative = positive 0-(-1) = 0 + 1 = 1

OpenStudy (anonymous):

thank you friend!

OpenStudy (welshfella):

yw